RC meeting and Run Plan, Tuesday October 25th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today to accommodate Hall C weekly meeting

  • Location: Counting house conference room

Useful Links

Current Status

Polarized Target

  • Laser work was preformed by Todd and and Jack. They found that 3 of the lasers are not working as expected.
  • Laser computer required Microsoft update and reboot (polhe1). Polhe1 and Polhe4 need to include exemptions so they do not automatically update and reboot.

BCM and Unser

  • Keith Cole and Dave Mack: When the beamline loses vacuum and is then pumped back down, BCM & Unser are detuned. Keith Cole could support a retuning of the Unser if there is an RF recovery
  • Dave Mack: For an asymmetry, the current BCM calibration would be fine. For cross sections.... may want another BCM calibration.

SBS GEM gas readout

  • Next access, will need to have Jack check the rPi

Other Topics

  • There is discussion for RV recovery on Thursday DAY (4 hours).
  • OPS and RF would like to study the ramp rate in Hall C while observing their ramp rate. This can be done in a dedicated beam studies, in which "parasitic beam" could be delivered to the halls. Jay anticipates the "parasitic beam" would be similar to what we are receiving now for production but likely with a higher trip rate. The test duration would be no more than 4 hours.
    • If it cannot be done due to Hall A's ramp rate (1 uA/s), they wonder if the ramp on the polarized helium-3 cell could be raised? To ~1.5 uA/s? Switch to a non-glass cell during this period?
  • Estimate of the total fraction of events collected?
    • Accumulated charge corrected for target polarization: 2.7 Coulombs
  • Excel spreadsheet created to track the DAQ crashes their sources was setup.
  • Subsystem updates?
  • A collaborator needs a hall walk through. Are there any others than may need one in order to coordinate with Jessie. If there is an RF recovery on Thursday, how long would be needed for the walk through? Are waivers available if they do not yet have it but have the other training?
  • Configuration and pass change moved to Monday (10/31/2022): [1]
    • Swapping of the leads to the power supply of the 3rd Moller quad
  • Opportunistic BCM calibration (higher current) before pass change?

Current Run Plan

For Monday swing and Tuesday OWL we will be focusing on polarized helium-3 production.

Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam delivery is interrupted for more than 4 hours, perform a checkout of the beam profile and raster follow Simona's procedure: [2]
  • Cross check the rate with what is expected: [3]
  • Cross check with the expected ion chamber readings: [4]

Notices in Effect for Shift Crew

Current Issues

  • Shift Openings: [5]

Planned Accesses to the Hall

Special Requests